# Angular integrals with three denominators via IBP, mass reduction, dimensional shift, and differential equations

### v1

@@ VII.1 Massless integral
Figure 2: Illustration of the initial condition and the integration path for the massless three-denominator angular integral depending on vectors $v_{1,2,3}$. The plot shows the slice $x_{0}=1$ of $d$-dimensional Minkowski space. The blue unit sphere depicts the intersection with the light-cone, where $v^{2}=0$. The symmetric configuration showcased by the solid arrows is used as the boundary value. In this configuration $I_{111}^{(0)}(1,1,1)=2\pi\mathrm{C}$, where $\mathrm{C}$ denotes Catalan’s constant. The dashed black lines showcase the integration path. First, we rotate $v_{1}$ along path $(a)$, which is in the direction of $\partial/\partial v_{12}$. The axis of rotation is along $v_{3}$. Second, we rotate $v_{1}$ along path $(b)$, which is in the direction of $\partial/\partial v_{13}$. The axis of rotation is along $v_{2}$. Third, we rotate $v_{2}$ along path $(c)$, which is in the direction of $\partial/\partial v_{23}$, reaching the final configuration with general $v_{12}$, $v_{13}$, $v_{23}$. The axis of rotation is the rotated $v_{1}$ shown in dashed red. Note that the volume of the tetrahedron spanned by the vectors $v_{1,2,3}$ in this subspace is proportional to $\sqrt{Y_{123}^{(0)}}$
@@
$J^{(0),d=6}_{111}(v_{12},v_{13},v_{23})=\int_{1}^{v_{12}}\!\!\!\!\mathrm{d}v\,j^{(0)}(v,v_{13},v_{23})+\int_{1}^{v_{13}}\!\!\!\!\mathrm{d}v\,j^{(0)}(v,1,v_{23})+\int_{1}^{v_{23}}\!\!\!\!\mathrm{d}v\,j^{(0)}(v,1,1)+J^{(0),d=6}_{111}(1,1,1)\,.$ (55)
We note that alternatively $J_{111}^{(0)}$ is also fixed by being the unique symmetric solution to eq. (53) with the boundary condition (54).
To evaluate the integrals from (55) in terms of (generalized) polylogarithms, the only remaining obstacle is the presence of the square root $\sqrt{Y_{123}^{(0)}}$. A general algorithmic approach to rationalize roots of this kind has been presented in [49]. In the present case the required change of variables is $v_{12}\rightarrow u$ with
@@
$Y_{123}^{(0)}=\frac{2u}{1+u^{2}}\,\sqrt{v_{13}(2-v_{13})v_{23}(2-v_{23})}\,.$ (58)
The integrals resulting from this change of variables take the form of $J_{0}(\delta)$ and $J_{2}(\beta,\delta)$ defined in appendix G. After considerable simplifications between the terms we obtain the result given in section VIII.1.
VII.2 Single-massive integral
To obtain the order $\varepsilon$ for $I_{111}^{(1)}$ in $d=4-2\varepsilon$ dimensions we only need the leading $\varepsilon^{0}$ term of eq. (52). Hence, we can again simply put $\varepsilon=0$ in this equation and plug in the known integrals with two and fewer denominators in $d=6$. Besides eq.(106) we use the dimensionally-shifted two-denominator integrals from appendix E leading to
$\frac{\partial J_{111}^{(1)}(-1)}{\partial v_{11}}=\frac{\pi}{\sqrt{Y_{123}^{(1)}}}$ $\left[\frac{(v_{12}-1)(v_{12}+v_{23}-v_{13}-v_{12}v_{23})\log\!\left(\frac{1+\sqrt{1-v_{11}}}{1-\sqrt{1-v_{11}}}\right)}{2(v_{12}(2-v_{12})-v_{11})\sqrt{1-v_{11}}}\right.$
@@ VII.2 Single-massive integral
Integrating the right-hand side starting at $v_{11}=0$ we can use the massless integral $J^{(0),d=6}_{111}$ as a boundary value,
$J_{111}^{(1),d=6}(v_{11})=J_{111}^{(0),d=6}+\int_{0}^{v_{11}}\mathrm{d}v\,j^{(1)}(v)\,.$ (60)
Note that when rescaling back from $J_{111}$ to $I_{111}$ the massless boundary term receives a $v_{11}$-dependent prefactor, namely
$I_{111}^{(1),d=6}=\frac{J_{111}^{(1),d=6}}{\sqrt{Y_{123}^{(1)}}}=\sqrt{\frac{Y_{123}^{(0)}}{Y_{123}^{(1)}}}\,I_{111}^{(0),d=6}+\frac{1}{\sqrt{Y_{123}^{(1)}}}\int_{0}^{v_{11}}\mathrm{d}v\,j^{(1)}(v)\,.$ (61)
@@
For the integrals additionally containing the square root $\sqrt{1-v_{11}}$, a change of variables $v_{11}\rightarrow s$ that simultaneously rationalizes both roots can be found using the approach of [49] and is given by
@@
$\sqrt{Y^{(1)}_{123}}=\frac{2s}{(1-s)(1+s)}\,\sqrt{v_{23}(2-v_{23})}\,\sqrt{1-w_{23}}\quad\text{and}\quad\sqrt{1-v_{11}}=\frac{1+s^{2}}{(1-s)(1+s)}\,\sqrt{1-w_{23}}\,.$ (65)
The integrals resulting from the change of variables take the form of $J_{0}(\delta)$ and $J_{1}(\alpha,\delta)$ defined in appendix G. After considerable simplifications between the terms we obtain the result given in section VIII.2.
VIII Results
In this section we present the $\varepsilon$-expansion for the massless, single-, double-, and triple-massive master integrals $I^{(m)}_{111}$ including order $\varepsilon$ terms. In combination with the recursion relations listed in appendix C these give the $\varepsilon$-expansion of all three-denominator angular integrals $I^{(m)}_{j_{1}j_{2}j_{3}}$ with integer-valued propagator powers $j_{i}$. The results are presented in a form that is manifestly real-valued for time- and light-like real vectors $v_{i}$ as will be the case in most applications. A detailed study of the non-trivial branch-cut structure, which arises outside this region, is beyond the scope of this paper.
VIII.1 Massless integral
@@
$I_{111}^{(0)}(v_{12},v_{13},v_{23};\varepsilon)=-\frac{\pi}{\varepsilon}\frac{(v_{12}+v_{13}+v_{23})}{v_{12}v_{13}v_{23}}$
@@
$+(v_{12}+v_{13}-v_{23})\,\mathrm{Li}_{2}\!\left(1-\frac{2}{v_{23}}\right)+\sqrt{Y_{123}^{(0)}}f^{(0)}_{Y}(v_{12},v_{13},v_{23})\right]+\mathcal{O}(\varepsilon^{2}),$ (66)
@@
$f^{(0)}_{Y}$ $(v_{12},v_{13},v_{23})=\frac{\sqrt{Y^{(0)}_{123}}}{\pi}I_{111}^{(0),d=6}=\frac{J_{111}^{(0)}(-1)}{\pi}$
$=\mathrm{Cl}_{2}\!\left(2\theta_{1}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{2}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{3}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{12}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{13}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{23}^{(0)}\right)-\mathrm{Cl}_{2}\!\left(2\theta_{123}^{(0)}\right)$ (67)
@@
$\theta_{i}^{(m)}=\arctan\frac{\sqrt{Y_{123}^{(m)}}}{v_{ij}+v_{ik}-v_{jk}-v_{ij}v_{ik}}\,,\quad\theta_{ij}^{(m)}=\arctan\frac{\sqrt{Y_{123}^{(m)}}}{v_{ik}+v_{jk}-v_{ij}}\,,\quad\theta_{ijk}^{(m)}=\arctan\frac{\sqrt{Y_{123}^{(m)}}}{4-v_{ij}-v_{ik}-v_{jk}}\,,$ (68)
where $i,j,k\in\{1,2,3\}$ always denote pairwise distinct indices. $\mathrm{Cl}_{2}$ is the Clausen function recalled in appendix H. Note that $f_{Y}^{(0)}(v_{12},v_{13},v_{23})$ is manifestly symmetric in its arguments. We recall our observation from section V that $\sqrt{Y_{123}^{(0)}}$ , which is given in eq. (39), is proportional to the volume of the tetrahedron spanned by $\vec{v}_{1,2,3}$ and hence vanishes if and only if the vectors $v_{1,2,3}$ become linearly dependent. In this case, the result reduces to a sum of two-denominator integrals matching the result one obtains by partial fractioning. The pole and finite parts are in agreement with the results from [4], which have been calculated by Mellin-Barnes techniques  [^4] Beware of the difference in normalization of the scalar products, to compare with [4] one needs to set $v_{ij}\rightarrow 2v_{ij}$ in our results..
@@
Further we write $\Omega_{123}$ for the solid angle covered by the vectors $(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$ – in other words, the area of the triangle on the unit sphere with edges at $(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$. Flipping one vector $\vec{v}_{i}\rightarrow-\vec{v}_{i}$, we define the additional solid angles $\Omega_{\bar{1}23}$, $\Omega_{1\bar{2}3}$ and, $\Omega_{12\bar{3}}$ where the flipped vector in each case is marked with a bar. Using Euler’s identity for spherical triangles from §23 of [50], given in modern notation in [51],
@@
Figure 3: Visualization of the geometrical quantities appearing in the result (72) for the massless integral in six dimensions. Analogous to figure 2, the picture shows the slice $x_{0}=1$ of six-dimensional Minkowski space, specifically the three-dimensional subspace defined by $\vec{v}_{1}$, $\vec{v}_{2}$, $\vec{v}_{3}$. These vectors span a tetrahedron with volume $\mathbb{V}(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$ shaded in gray. Note that the factor of $6$ in eq. (72) is not from $d=6$ but originates from eq. (34). The dihedral angle between the faces that meet at $\vec{v}_{i}$ is denoted by $\phi_{i}$, i.e. e.g. $\phi_{1}$ is the angle between the face spanned by $(\vec{v}_{1},\vec{v}_{2})$ and $(\vec{v}_{1},\vec{v}_{3})$. They contribute with $\mathrm{Cl}_{2}(2\phi_{i})$ in eq. (72). The $\Omega$ are solid angles, i.e. areas of triangles on the the unit sphere. The red area $\Omega_{\bar{1}23}$ has corners at $(-\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$, the yellow area $\Omega_{1\bar{2}3}$ has corners at $(\vec{v}_{1},-\vec{v}_{2},\vec{v}_{3})$, and the blue area $\Omega_{\bar{1}23}$ has corners at $(\vec{v}_{1},\vec{v}_{2},-\vec{v}_{3})$. They contribute with $\mathrm{Cl}_{2}(\Omega)$ in eq. (72). The black area $\Omega_{123}$ is the solid angle covered by $(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$ and contributes negatively with $-\mathrm{Cl}_{2}(\Omega)$ in eq. (72).
@@
$I_{111}^{(0),d=6}=\frac{\pi}{6\mathbb{V}(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})}\Big{[}\mathrm{Cl}_{2}(2\phi_{1})+\mathrm{Cl}_{2}(2\phi_{2})+\mathrm{Cl}_{2}(2\phi_{3})+\mathrm{Cl}_{2}(\Omega_{\bar{1}23})+\mathrm{Cl}_{2}(\Omega_{1\bar{2}3})+\mathrm{Cl}_{2}(\Omega_{12\bar{3}})-\mathrm{Cl}_{2}(\Omega_{123})\Big{]},$ (72)
@@
VIII.2 Single-massive integral
@@
$I_{111}^{(1)}(v_{12},v_{13},v_{23},v_{11};\varepsilon)=-\frac{\pi}{\varepsilon}\frac{(v_{12}+v_{13})}{v_{12}v_{13}v_{23}}$
@@
$\qquad+v_{13}(v_{11}v_{23}+v_{12}^{2}-v_{12}v_{13}-v_{12}v_{23})\left(\mathrm{Li}_{2}\!\left(1-\frac{v_{12}}{1+\sqrt{1-v_{11}}}\right)+\mathrm{Li}_{2}\!\left(1-\frac{v_{12}}{1-\sqrt{1-v_{11}}}\right)+\frac{1}{4}\log\frac{v_{11}}{v_{12}^{2}}\right)$
$\qquad+v_{12}(v_{11}v_{23}+v_{13}^{2}-v_{12}v_{13}-v_{13}v_{23})\left(\mathrm{Li}_{2}\!\left(1-\frac{v_{13}}{1+\sqrt{1-v_{11}}}\right)+\mathrm{Li}_{2}\!\left(1-\frac{v_{13}}{1-\sqrt{1-v_{11}}}\right)+\frac{1}{4}\log\frac{v_{11}}{v_{13}^{2}}\right)$
$\qquad+v_{12}v_{13}\sqrt{Y_{123}^{(1)}}f_{Y}^{(1)}(v_{12},v_{13},v_{23},v_{11})\left.\vphantom{\frac{}{}}\right],$ (73)
@@
$f_{Y}^{(1)}$ $(v_{12},v_{13},v_{23},v_{11})=\frac{\sqrt{Y_{123}^{(1)}}}{\pi}I_{111}^{(1),d=6}=\frac{J_{111}^{(1)}(-1)}{\pi}=f_{Y}^{(0)}+\int_{0}^{v_{11}}\mathrm{d}v\,j^{(1)}(v)$
$=$ $\,\mathrm{Cl}_{2}\!\left(2\theta_{1}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{2}^{(1)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{3}^{(1)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{12}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{13}^{(0)}\right)+2\mathrm{Cl}_{2}\!\left(2\theta_{23}^{(1)}\right)-\mathrm{Cl}_{2}\!\left(2\theta_{23}^{(0)}\right)-\mathrm{Cl}_{2}\!\left(2\theta_{123}^{(0)}\right)$
@@
$+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}+2\phi_{12,+}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}-2\phi_{12,+}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}+2\phi_{12,-}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}-2\phi_{12,-}^{(0)}\right)$
$-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}+2\phi_{12,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}-2\phi_{12,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}+2\phi_{2,+}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}-2\phi_{12,+}^{(0)}\right)$
@@
$+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(0)}+2\phi_{13,+}^{(1)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(1)}-2\phi_{13,+}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}+2\phi_{13,-}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}-2\phi_{13,-}^{(0)}\right)$
$-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(1)}+2\phi_{13,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(1)}-2\phi_{13,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}+2\phi_{13,+}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}-2\phi_{13,+}^{(0)}\right)$

### v4

@@ 7.1 Massless integral
Figure 2: Illustration of the initial condition and the integration path for the massless three-denominator angular integral depending on vectors $v_{1,2,3}$. The plot shows the slice $x_{0}=1$ of $d$-dimensional Minkowski space. The blue unit sphere depicts the intersection with the light-cone, where $v^{2}=0$. The symmetric configuration showcased by the solid arrows is used as the boundary value. In this configuration $I_{1,1,1}^{(0)}(1,1,1)=2\pi\mathrm{C}$, where $\mathrm{C}$ denotes Catalan’s constant. The dashed black lines showcase the integration path. First, we rotate $v_{1}$ along path $(a)$, which is in the direction of $\partial/\partial v_{12}$. The axis of rotation is along $v_{3}$. Second, we rotate $v_{1}$ along path $(b)$, which is in the direction of $\partial/\partial v_{13}$. The axis of rotation is along $v_{2}$. Third, we rotate $v_{2}$ along path $(c)$, which is in the direction of $\partial/\partial v_{23}$, reaching the final configuration with general $v_{12}$, $v_{13}$, $v_{23}$. The axis of rotation is the rotated $v_{1}$ shown in dashed red. Note that the volume of the tetrahedron spanned by the vectors $v_{1,2,3}$ in this subspace is proportional to $\sqrt{Y_{123}^{(0)}}$ .
@@
$J^{(0),d=6}_{1,1,1}(v_{12},v_{13},v_{23})=$ $\int_{1}^{v_{12}}\!\!\!\!\mathrm{d}v\,j^{(0)}(v,v_{13},v_{23})+\int_{1}^{v_{13}}\!\!\!\!\mathrm{d}v\,j^{(0)}(v,1,v_{23})$
$+$ $\int_{1}^{v_{23}}\!\!\!\!\mathrm{d}v\,j^{(0)}(v,1,1)+J^{(0),d=6}_{1,1,1}(1,1,1)\,.$ (56)
We note that alternatively $J_{1,1,1}^{(0)}$ is also fixed by being the unique symmetric solution to eq. (54) with the boundary condition (55).
To evaluate the integrals from (56) in terms of (generalized) polylogarithms, the only remaining obstacle is the presence of the square root $\sqrt{Y_{123}^{(0)}}$. A general algorithmic approach to rationalize roots of this kind has been presented in Besier:2018 . In the present case the required change of variables is $v_{12}\rightarrow u$ with
@@
$\sqrt{Y_{123}^{(0)}}=\frac{2u}{1+u^{2}}\,\sqrt{v_{13}(2-v_{13})v_{23}(2-v_{23})}\,.$ (59)
The integrals resulting from this change of variables take the form of $J_{0}(\delta)$ and $J_{2}(\beta,\delta)$ defined in appendix G. After considerable simplifications between the terms we obtain the result given in section 8.1.
7.2 Single-massive integral
To obtain the order $\varepsilon$ for $I_{1,1,1}^{(1)}$ in $d=4-2\varepsilon$ dimensions we only need the leading $\varepsilon^{0}$ term of eq. (53). Hence, we can again simply put $\varepsilon=0$ in this equation and plug in the known integrals with two and fewer denominators in $d=6$. Besides eq.(107) we use the dimensionally-shifted two-denominator integrals from appendix E leading to
$\frac{\partial J_{1,1,1}^{(1)}(-1)}{\partial v_{11}}=\frac{\pi}{\sqrt{Y_{123}^{(1)}}}\left[\frac{(v_{12}-1)(v_{12}+v_{23}-v_{13}-v_{12}v_{23})\log\!\left(\frac{1+\sqrt{1-v_{11}}}{1-\sqrt{1-v_{11}}}\right)}{2(v_{12}(2-v_{12})-v_{11})\sqrt{1-v_{11}}}\right$.
@@ 7.2 Single-massive integral
Integrating the right-hand side starting at $v_{11}=0$ we can use the massless integral $J^{(0),d=6}_{1,1,1}$ as a boundary value,
$J_{1,1,1}^{(1),d=6}(v_{11})=J_{1,1,1}^{(0),d=6}+\int_{0}^{v_{11}}\mathrm{d}v\,j^{(1)}(v)\,.$ (61)
Note that when rescaling back from $J_{1,1,1}$ to $I_{1,1,1}$ the massless boundary term receives a $v_{11}$-dependent prefactor, namely
$I_{1,1,1}^{(1),d=6}=\frac{J_{1,1,1}^{(1),d=6}}{\sqrt{Y_{123}^{(1)}}}=\sqrt{\frac{Y_{123}^{(0)}}{Y_{123}^{(1)}}}\,I_{1,1,1}^{(0),d=6}+\frac{1}{\sqrt{Y_{123}^{(1)}}}\int_{0}^{v_{11}}\mathrm{d}v\,j^{(1)}(v)\,.$ (62)
@@
For the integrals additionally containing the square root $\sqrt{1-v_{11}}$, a change of variables $v_{11}\rightarrow s$ that simultaneously rationalizes both roots can be found using the approach of Besier:2018 and is given by
@@
$\sqrt{Y^{(1)}_{123}}=\frac{2s}{(1-s)(1+s)}\,\sqrt{v_{23}(2-v_{23})}\,\sqrt{1-w_{23}}\;\;\text{and}\;\;\sqrt{1-v_{11}}=\frac{1+s^{2}}{(1-s)(1+s)}\,\sqrt{1-w_{23}}\,.$ (66)
The integrals resulting from the change of variables take the form of $J_{0}(\delta)$ and $J_{1}(\alpha,\delta)$ defined in appendix G. After considerable simplifications between the terms we obtain the result given in section 8.2.
8 Results
In this section we present the $\varepsilon$-expansion for the massless, single-, double-, and triple-massive master integrals $I^{(m)}_{1,1,1}$ including order $\varepsilon$ terms. In combination with the recursion relations listed in appendix C these give the $\varepsilon$-expansion of all three-denominator angular integrals $I^{(m)}_{j_{1},j_{2},j_{3}}$ with integer-valued propagator powers $j_{i}$. The results are presented in a form that is manifestly real-valued for time- and light-like real vectors $v_{i}$ as will be the case in most applications. A detailed study of the non-trivial branch-cut structure, which arises outside this region, is beyond the scope of this paper.
8.1 Massless integral
@@
$I_{1,1,1}^{(0)}(v_{12},v_{13},v_{23};\varepsilon)=-\frac{\pi}{\varepsilon}\frac{(v_{12}+v_{13}+v_{23})}{v_{12}v_{13}v_{23}}$
@@
[…]
@@
$f^{(0)}_{Y}(v_{12},v_{13},v_{23})$ $=\frac{\sqrt{Y^{(0)}_{123}}}{\pi}I_{1,1,1}^{(0),d=6}=\frac{J_{1,1,1}^{(0)}(-1)}{\pi}$
$=\mathrm{Cl}_{2}\!\left(2\theta_{1}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{2}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{3}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{12}^{(0)}\right)$
$+\mathrm{Cl}_{2}\!\left(2\theta_{13}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{23}^{(0)}\right)-\mathrm{Cl}_{2}\!\left(2\theta_{123}^{(0)}\right)$ (68)
@@
$\theta_{i}^{(m)}=\arctan\frac{\sqrt{Y_{123}^{(m)}}}{v_{ij}+v_{ik}-v_{jk}-v_{ij}v_{ik}}\,,\quad\theta_{ij}^{(m)}=\arctan\frac{\sqrt{Y_{123}^{(m)}}}{v_{ik}+v_{jk}-v_{ij}}\,$,
$\theta_{ijk}^{(m)}=\arctan\frac{\sqrt{Y_{123}^{(m)}}}{4-v_{ij}-v_{ik}-v_{jk}}\,$, (69)
where $i,j,k\in\{1,2,3\}$ always denote pairwise distinct indices. $\mathrm{Cl}_{2}$ is the Clausen function recalled in appendix H. Note that $f_{Y}^{(0)}(v_{12},v_{13},v_{23})$ is manifestly symmetric in its arguments. We recall our observation from section 5 that $\sqrt{Y_{123}^{(0)}}$ , which is given in eq. (40), is proportional to the volume of the tetrahedron spanned by $\vec{v}_{1,2,3}$ and hence vanishes if and only if the vectors $v_{1,2,3}$ become linearly dependent. In this case, the result reduces to a sum of two-denominator integrals matching the result one obtains by partial fractioning. The pole and finite parts are in agreement with the results from Somogyi:2011 , which have been calculated by Mellin-Barnes techniques  [^4] Beware of the difference in normalization of the scalar products, to compare with Somogyi:2011 one needs to set $v_{ij}\rightarrow 2v_{ij}$ in our results..
Figure 3: Visualization of the geometrical quantities appearing in the result (73) for the massless integral in six dimensions. Analogous to figure 2, the picture shows the slice $x_{0}=1$ of six-dimensional Minkowski space, specifically the three-dimensional subspace defined by $\vec{v}_{1}$, $\vec{v}_{2}$, $\vec{v}_{3}$. These vectors span a tetrahedron with volume $\mathbb{V}(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$ shaded in gray. Note that the factor of $6$ in eq. (73) is not from $d=6$ but originates from eq. (35). The dihedral angle between the faces that meet at $\vec{v}_{i}$ is denoted by $\phi_{i}$, i.e. e.g. $\phi_{1}$ is the angle between the face spanned by $(\vec{v}_{1},\vec{v}_{2})$ and $(\vec{v}_{1},\vec{v}_{3})$. They contribute with $\mathrm{Cl}_{2}(2\phi_{i})$ in eq. (73). The $\Omega$ are solid angles, i.e. areas of triangles on the the unit sphere. The red area $\Omega_{\bar{1}23}$ has corners at $(-\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$, the yellow area $\Omega_{1\bar{2}3}$ has corners at $(\vec{v}_{1},-\vec{v}_{2},\vec{v}_{3})$, and the blue area $\Omega_{\bar{1}23}$ has corners at $(\vec{v}_{1},\vec{v}_{2},-\vec{v}_{3})$. They contribute with $\mathrm{Cl}_{2}(\Omega)$ in eq. (73). The black area $\Omega_{123}$ is the solid angle covered by $(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$ and contributes negatively with $\mathrm{Cl}_{2}(\Omega)$ in eq. (73).
@@
Further we write $\Omega_{123}$ for the solid angle covered by the vectors $(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$ – in other words, the area of the triangle on the unit sphere with edges at $(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})$. Flipping one vector $\vec{v}_{i}\rightarrow-\vec{v}_{i}$, we define the additional solid angles $\Omega_{\bar{1}23}$, $\Omega_{1\bar{2}3}$ and, $\Omega_{12\bar{3}}$ where the flipped vector in each case is marked with a bar. Using Euler’s identity for spherical triangles from §23 of Euler:1781 , given in modern notation in Eriksson:1990 ,
@@
$I_{1,1,1}^{(0),d=6}=\frac{\pi}{6\mathbb{V}(\vec{v}_{1},\vec{v}_{2},\vec{v}_{3})}$ $\Big[\mathrm{Cl}_{2}(2\phi_{1})+\mathrm{Cl}_{2}(2\phi_{2})+\mathrm{Cl}_{2}(2\phi_{3})+\mathrm{Cl}_{2}(\Omega_{\bar{1}23})$
$+\mathrm{Cl}_{2}(\Omega_{1\bar{2}3})+\mathrm{Cl}_{2}(\Omega_{12\bar{3}})-\mathrm{Cl}_{2}(\Omega_{123})\Big],$ (73)
@@
8.2 Single-massive integral
@@
$I_{1,1,1}^{(1)}(v_{12},v_{13},v_{23},v_{11};\varepsilon)=-\frac{\pi}{\varepsilon}\frac{(v_{12}+v_{13})}{v_{12}v_{13}v_{23}}$
@@
$\qquad+v_{13}(v_{11}v_{23}+v_{12}^{2}-v_{12}v_{13}-v_{12}v_{23})$
$\qquad\quad\times\left(\mathrm{Li}_{2}\!\left(1-\frac{v_{12}}{1+\sqrt{1-v_{11}}}\right)+\mathrm{Li}_{2}\!\left(1-\frac{v_{12}}{1-\sqrt{1-v_{11}}}\right)+\frac{1}{4}\log\frac{v_{11}}{v_{12}^{2}}\right)$
$\qquad+v_{12}(v_{11}v_{23}+v_{13}^{2}-v_{12}v_{13}-v_{13}v_{23})$
$\qquad\quad\times\left(\mathrm{Li}_{2}\!\left(1-\frac{v_{13}}{1+\sqrt{1-v_{11}}}\right)+\mathrm{Li}_{2}\!\left(1-\frac{v_{13}}{1-\sqrt{1-v_{11}}}\right)+\frac{1}{4}\log\frac{v_{11}}{v_{13}^{2}}\right)$
[…]
@@
$f_{Y}^{(1)}$ $(v_{12},v_{13},v_{23},v_{11})=\frac{\sqrt{Y_{123}^{(1)}}}{\pi}I_{1,1,1}^{(1),d=6}=\frac{J_{1,1,1}^{(1)}(-1)}{\pi}=f_{Y}^{(0)}+\int_{0}^{v_{11}}\mathrm{d}v\,j^{(1)}(v)$
$=$ $\,\mathrm{Cl}_{2}\!\left(2\theta_{1}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{2}^{(1)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{3}^{(1)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{12}^{(0)}\right)+\mathrm{Cl}_{2}\!\left(2\theta_{13}^{(0)}\right)$
$+2\mathrm{Cl}_{2}\!\left(2\theta_{23}^{(1)}\right)-\mathrm{Cl}_{2}\!\left(2\theta_{23}^{(0)}\right)-\mathrm{Cl}_{2}\!\left(2\theta_{123}^{(0)}\right)$
@@
$+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}+2\phi_{12,+}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}-2\phi_{12,+}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}+2\phi_{12,-}^{(0)}\right)$
$+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}-2\phi_{12,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}+2\phi_{12,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,+}^{(1)}-2\phi_{12,-}^{(0)}\right)$
$-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}+2\phi_{2,+}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{12,-}^{(1)}-2\phi_{12,+}^{(0)}\right)$
@@
$+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(0)}+2\phi_{13,+}^{(1)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(1)}-2\phi_{13,+}^{(0)}\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}+2\phi_{13,-}^{(0)}\right)$
$+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}-2\phi_{13,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(1)}+2\phi_{13,-}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,+}^{(1)}-2\phi_{13,-}^{(0)}\right)$
$-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}+2\phi_{13,+}^{(0)}\right)-\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\phi_{13,-}^{(1)}-2\phi_{13,+}^{(0)}\right)$